# Rolled Gaussian process models for curves on manifolds

### v1

@@ 7 Application: orientation of end-effector of a robot arm, curves on $SO(3)$
The data are orientations of the end-effector of a Franka robot arm as it was guided $n=60$ times by the third author to perform a task to deposit the contents of a dustpan into a bin. The orientation can be represented by an element of ${SO}(3)$, the set of $3$-by-$3$ orthogonal matrices with unit determinant, that describes its rotation from an initial reference orientation. An element of $SO(3)$ may be parameterised as an unsigned unit quaternion, and $SO(3)$ is hence identified with $\mathbb{S}^{3}$ modulo the antipodal map (Prentice, 1986). Fixing the sign of each data point (e.g. to positive) then identifies $SO(3)$ with a hemisphere of $\mathbb{S}^{3}$, which is our manifold $M$ equipped with the induced geometry of $\mathbb{S}^{3}$, whose geometry is inherited from $\mathbb{R}^{4}$.
The curves, $X_{i}$, are recorded at $r=100$ time points. We choose basis dimension $k=10$, and set $b=e_{1}$ and $U=(e_{2},e_{3},e_{4})$, where $\{e_{i}\}$ is the standard basis for $\mathbb{R}^{4}$. We compute the estimated mean curve $\hat{\Gamma}=(b1_{r}^{\top}+U\hat{M}_{w}\Phi)_{b}^{\uparrow}$, with $\hat{M}_{w}$ estimated as in (4.6). Then the unwrapping coordinates, $H_{i}=H(X_{i};\hat{\Gamma})$ for $H(\cdot;\cdot)$ in (4.2), are curves in $\mathbb{R}^{3}$; these are shown as blue curves in Fig. 4(a), together with corresponding unrolled mean, $H(\hat{\Gamma};\hat{\Gamma})$, plotted in red. The curves have a common starting point at $t=0$, shown near the top-left in this plot; for small $t$ the variability is small, but then at larger $t$ there is much larger variability, especially following a kink point that corresponds to the dustpan being turned to empty its contents. We further estimate the covariance matrices $U_{w}$ and $V_{w}$ in (4.1) using (4.8, 4.9). As a visual appraisal of the fitted model, Fig. 4(b) shows $n=60$ realisations from $\mathcal{RMN}(\hat{M}_{w},\hat{U}_{w},\hat{V}_{w};b,U)$, using the same unwrapping coordinates and projection as in Fig. 4(a). These simulated curves are smoother than the real data, which is a consequence of the basis used, but they have similar heteroscedastic variation.
@@ 8 Discussion
The rolling approach presented here extends beyond the Gaussian process setting. Future work will explore rolling and wrapping of other processes on $\mathbb{R}^{d}$ to develop other useful new models on manifolds $M$.
For our examples, we leveraged closed-form expressions for exponential, inverse exponential, and parallel transport maps when computing with the four maps in §3.1. On manifolds lacking these closed-form expressions, approximations via retractions and their inverses offer practical alternatives (Absil et al., 2009).
Within the Gaussian process setting of this paper, we have specialised to a particular parameterization of the mean and covariance, and assumed common and equispace observation times; these provide modelling and computational convenience but are straightforward to relax. A further generalization is to incorporate a model of nuisance variation in the time variable via warping. This would accommodate the possibility of time registration of observed curves, often valuable in analogous Euclidean settings (Zhang et al., 2018; Su et al., 2014). However, for rolled models, warping combined with rolling raises challenges for computation and inference that remain to be resolved. The foregoing points present exciting opportunities to exploit the strategy of rolling in wider settings involving data on manifolds.
@@ 9 Acknowledgements
We thank Huiling Le for helpful discussions. SP acknowledges support from a grant from the Engineering and Physical Sciences Research Council (EPSRC EP/T003928/1). KB acknowledges support from grants from the Engineering and Physical Sciences Research Council (EPSRC EP/X022617/1), National Science Foundation (NSF 2015374), and the National Insitutes of Health (NIH R37-CA21495).
10 Appendix
10.1 Gaussian distribution on inner product spaces
@@
With respect to $(\Omega,\mathcal{F},\mathbb{P})$ suppose $X_{1}$ and $X_{2}$ are two random vectors on $(X_{1},\langle\cdot,\cdot\rangle_{X_{1}})$ and $(X_{2},\langle\cdot,\cdot\rangle_{X_{2}})$, respectively, with covariances $\Sigma_{1}$ and $\Sigma_{2}$. The random vector $(X_{1},X_{2})$ then assumes values in the direct sum space $X_{1}\oplus X_{2}$ equipped with the inner product $\langle\cdot,\cdot\rangle_{X_{1}}+\langle\cdot,\cdot\rangle_{X_{2}}$. A well-defined positive semidefinite bilinear map $(x_{1},x_{2})\mapsto\text{Cov}(\langle x_{1},X_{1}\rangle_{X_{1}},\langle x_{2},X_{2}\rangle_{X_{2}})$ ensures the existence of a unique operator $\Sigma_{12}\in\mathcal{L}(X_{1},X_{2})$. Then, the linear operator $\Sigma\in\mathcal{L}(X_{1}\oplus X_{2},X_{1}\oplus X_{2})$ defined as $\Sigma(x_{1},x_{2})=(\Sigma_{11}x_{1}+\Sigma_{12}x_{2},\Sigma_{12}^{\top}x_{1},\Sigma_{22}x_{2})$, is known as the covariance of the random vector $(X_{1},X_{2})$ in $X_{1}\oplus X_{2}$. The above definition of extends easily to random vectors $(X_{1},\ldots,X_{m})$ with values in $X_{1}\oplus\cdots\oplus X_{m}$.
10.2 Orthonormal frame along a curve $\gamma$
@@
10.3 Computations with discrete representations of curves
@@
$=\gamma^{\downarrow}(t_{j})+\int_{t_{j}}^{t}P^{\tilde{\gamma}}_{t_{0}\leftarrow t_{j}}P^{\tilde{\gamma}}_{t_{j}\leftarrow s}\dot{\tilde{\gamma}}(s)\mathrm{d}s=\gamma^{\downarrow}(t_{j})+P^{\tilde{\gamma}}_{t_{0}\leftarrow t_{j}}\exp^{-1}_{\gamma(t_{j})}{\tilde{\gamma}}(t),$ (10.1)
using recursion, and using that along the geodesic segment between ${\tilde{\gamma}}(t_{j})$ and ${\tilde{\gamma}}(t)$, $\int_{t_{j}}^{t}P^{\tilde{\gamma}}_{t_{j}\leftarrow s}\dot{\tilde{\gamma}}(s)\mathrm{d}s=t\dot{\tilde{\gamma}}(t_{j})=\exp^{-1}_{\gamma(t_{j})}\tilde{\gamma}(t).$ Hence, via equation (10.1), $\Gamma^{\downarrow}=\{\gamma^{\downarrow}(t_{1}),\ldots,\gamma^{\downarrow}(t_{m})\}$ can be determined from $\Gamma=\{\gamma(t_{1}),\ldots,\gamma(t_{m})\}$. Similarly, via the rearrangement of (10.1) as
$\gamma(t)=\exp_{\gamma(t_{j})}\left[P^{\tilde{\gamma}}_{t_{j}\leftarrow t_{0}}\left\{\gamma^{\downarrow}(t)-\gamma^{\downarrow}(t_{j})\right\}\right],$ (10.2)
@@
10.4 Expressions for manifold $\mathbb{S}^{d}$
The embedded representation of the manifold is $M=\mathbb{S}^{d}=\{x\in\mathbb{R}^{q}:x^{\top}x=1\}$ for $q=d+1$. Suppose that $p,p^{\prime},x\in M$; $y\in T_{p}M$; and $c$ is a geodesic curve connecting $p$ to $p^{\prime}\in M\backslash\mathcal{C}(p)$ such that $c(0)=p$ and $c(1)=p^{\prime}$. Then (Absil et al., 2009, for example),
$\exp_{p}^{-1}(x)$ $=u\{x-p\cos(u)\}/\sin(u)\in T_{p}M,$
$\exp_{p}(y)$ $=p\cos(\|y\|)+(y/\|y\|)\sin(y/\|y\|)\in M,$
$P_{1\leftarrow 0}^{c}y$ $=\left[I_{d}+\{\cos(v)-1\}ww^{\top}-\sin(v)pw^{\top}\right]y\in T_{p^{\prime}}M,$
@@ 10.4 Expressions for manifold $\mathbb{S}^{d}$
10.5 Expressions for manifold $\operatorname{Sym}_{>0}(d)$.
@@
$\exp_{P}^{-1}(X)$ $=P^{1/2}\log\left(P^{-1/2}XP^{-1/2}\right)P^{1/2}\in T_{P}M,$
$\exp_{P}(Y)$ $=P^{1/2}\exp\left(P^{-1/2}YP^{-1/2}\right)P^{1/2}\in M,$

### v2

@@ 9 Application: orientation of end-effector of a robot arm, curves on $SO(3)$
The data are orientations of the end-effector of a Franka robot arm as it was guided $n=60$ times by the third author to perform a task to deposit the contents of a dustpan into a bin. The orientation can be represented by an element of SO(3){SO}(3), the set of 33-by-33 orthogonal matrices with unit determinant, that describes its rotation from an initial reference orientation. An element of $SO(3)$ may be parameterised as an unsigned unit quaternion, and $SO(3)$ is hence identified with $\mathbb{S}^{3}$ modulo the antipodal map (Prentice, 1986). Fixing the sign of each data point then identifies $SO(3)$ with a hemisphere of $\mathbb{S}^{3}$, which is our manifold $M$ equipped with the induced geometry of $\mathbb{S}^{3}$, whose geometry is inherited from $\mathbb{R}^{4}$.
[…]
@@ 10 Discussion
Sample path and distributional properties of Euclidean Gaussian processes have been extensively studied. There is much to investigate about analogous properties of the rolled Gaussian process, and this constitutes ongoing work. Moreover, the rolling approach presented here extends beyond the Gaussian process setting, and in future work we will explore such models.
For our examples, we leveraged closed-form expressions for exponential, inverse exponential, and parallel transport maps when computing with the four maps in §3. On manifolds lacking these closed-form expressions, approximations via retractions and their inverses offer practical alternatives (Absil et al., 2009). See also Nguyen and Sommer (2025) for matrix manifolds.
Within the Gaussian process setting of this paper, we have specialised to a particular parameterization of the mean and covariance, and assumed common and equispace observation times; these provide modelling and computational convenience but are straightforward to relax. When sample curves are observed sparsely at irregularly spaced observation times, and perhaps in the presence of measurement error, some form of smoothing or resampling of the curves is required to ensure that the design points are common to all curves. Once this is done, the proposed estimators of the mean $m$ and $K$ covariance functions of the Euclidean Gaussian process, defined using unwrapping coordinates §6.1 may be used. In future work, we will develop the corresponding asymptotic theory for such estimators.
Relatedly, a further generalization is to incorporate a model of nuisance variation in the time variable via warping. This would accommodate the possibility of time registration of observed curves, often valuable in analogous Euclidean settings (Zhang et al., 2018a; Su et al., 2014). However, for rolled models, warping combined with rolling raises challenges for computation and inference that remain to be resolved.
@@ 11 Acknowledgements
We thank Huiling Le for helpful discussions. SP acknowledges support from a grant from the Engineering and Physical Sciences Research Council (EPSRC EP/T003928/1). KB acknowledges support from grants from the Engineering and Physical Sciences Research Council (EP/Z003377/1; EPSRC EP/X022617/1), National Science Foundation (NSF 2015374), and the National Insitutes of Health (NIH R37-CA21495).
Appendix A Appendix
A.1 Gaussian distribution on inner product spaces
@@
With respect to $(\Omega,\mathcal{F},\mathbb{P})$ suppose $X_{1}$ and $X_{2}$ are two random vectors on $(X_{1},\langle\cdot,\cdot\rangle_{X_{1}})$ and $(X_{2},\langle\cdot,\cdot\rangle_{X_{2}})$, respectively, with covariances $\Sigma_{1}$ and $\Sigma_{2}$. The random vector $(X_{1},X_{2})$ then assumes values in the direct sum space $X_{1}\oplus X_{2}$ equipped with the inner product $\langle\cdot,\cdot\rangle_{X_{1}}+\langle\cdot,\cdot\rangle_{X_{2}}$. A well-defined bilinear map $(x_{1},x_{2})\mapsto\text{Cov}(\langle x_{1},X_{1}\rangle_{X_{1}},\langle x_{2},X_{2}\rangle_{X_{2}})$ that is positive semidefinite ensures the existence of a unique operator $\Sigma_{12}\in\mathcal{L}(X_{1},X_{2})$. Then, the linear operator $\Sigma\in\mathcal{L}(X_{1}\oplus X_{2},X_{1}\oplus X_{2})$ defined as $\Sigma(x_{1},x_{2})=(\Sigma_{11}x_{1}+\Sigma_{12}x_{2},\Sigma_{12}^{\top}x_{1},\Sigma_{22}x_{2})$, is known as the covariance of the random vector $(X_{1},X_{2})$ in $X_{1}\oplus X_{2}$. The above definition of extends easily to random vectors $(X_{1},\ldots,X_{m})$ with values in $X_{1}\oplus\cdots\oplus X_{m}$.
A.2 Orthonormal frame along a curve $\gamma$
@@
A.3 $C^{k}$ curves in $M$
Given a chart $(U,\psi)$ with open $U\subset M$ and smooth $\psi:U\to V$, where $V$ is an open subset of $\mathbb{R}^{d}$, the coordinate representative of a curve $c:[0,1]\to M$ is the Euclidean curve $\tilde{c}=\psi\circ c$. Then $c$ is a $C^{1}$ curve if $\tilde{c}$ is $C^{k}$ with respect to the norm $\|\tilde{c}\|_{C^{k}}:=\sum_{j=0}^{k}\sup_{t}\|D^{j}\tilde{c}(t)\|$, where $D^{j}$ is the $j$th derivative operator with $D^{0}$ equalling the identity. Then, a sequence of curves in $M$ converge in the weak $C^{k}$ topology to a limit if their coordinate representatives converge to the corresponding representative for the limit with respect to the $C^{k}$ norm (Chapter 2 Hirsch, 2012).
Appendix B Numerical implementation
B.1 Computations with discrete representations of curves
@@
$=\gamma^{\downarrow}(t_{j})+\int_{t_{j}}^{t}P^{\tilde{\gamma}}_{t_{0}\leftarrow t_{j}}P^{\tilde{\gamma}}_{t_{j}\leftarrow s}\dot{\tilde{\gamma}}(s)\mathrm{d}s=\gamma^{\downarrow}(t_{j})+P^{\tilde{\gamma}}_{t_{0}\leftarrow t_{j}}\exp^{-1}_{\gamma(t_{j})}{\tilde{\gamma}}(t),$ (B.1)
using recursion, and using that along the geodesic segment between ${\tilde{\gamma}}(t_{j})$ and ${\tilde{\gamma}}(t)$, $\int_{t_{j}}^{t}P^{\tilde{\gamma}}_{t_{j}\leftarrow s}\dot{\tilde{\gamma}}(s)\mathrm{d}s=t\dot{\tilde{\gamma}}(t_{j})=\exp^{-1}_{\gamma(t_{j})}\tilde{\gamma}(t).$ Hence, via equation (B.1), $\Gamma^{\downarrow}=\{\gamma^{\downarrow}(t_{1}),\ldots,\gamma^{\downarrow}(t_{m})\}$ can be determined from $\Gamma=\{\gamma(t_{1}),\ldots,\gamma(t_{m})\}$. Similarly, via the rearrangement of (B.1) as
$\gamma(t)=\exp_{\gamma(t_{j})}\left[P^{\tilde{\gamma}}_{t_{j}\leftarrow t_{0}}\left\{\gamma^{\downarrow}(t)-\gamma^{\downarrow}(t_{j})\right\}\right],$ (B.2)
@@
B.2 Expressions for manifold $\mathbb{S}^{d}$
The embedded representation of the manifold is $\mathbb{S}^{d}=\{x\in\mathbb{R}^{q}:x^{\top}x=1\}$ for $q=d+1$. Suppose that $p,p^{\prime},x\in M$; $y\in T_{p}M$; and $c$ is a geodesic curve connecting $p$ to $p^{\prime}\in M\backslash\mathcal{C}(p)$ such that $c(0)=p$ and $c(1)=p^{\prime}$. Then (Absil et al., 2009, for example),
$\exp_{p}^{-1}(x)$ $=u\{x-p\cos(u)\}/\sin(u)\in T_{p}M,\quad x\notin\mathcal{C}(p);$
$\exp_{p}(y)$ $=p\cos(\|y\|)+(y/\|y\|)\sin(y/\|y\|)\in M;$
$P_{1\leftarrow 0}^{c}y$ $=\left[I_{d}+\{\cos(v)-1\}ww^{\top}-\sin(v)pw^{\top}\right]y\in T_{p^{\prime}}M;$
@@ B.2 Expressions for manifold $\mathbb{S}^{d}$
B.3 Expressions for manifold $\operatorname{Sym}_{>0}(d)$.
@@
$\exp_{P}^{-1}(X)$ $=P^{1/2}\log\left(P^{-1/2}XP^{-1/2}\right)P^{1/2}\in T_{P}M;$
$\exp_{P}(Y)$ $=P^{1/2}\exp\left(P^{-1/2}YP^{-1/2}\right)P^{1/2}\in M;$